Solve each system of equations 4x-y=-5 4x-3y=8
-2x+1/2y=1 2x+y=14

Answer by mananth(16949)

1 x + 1 y = -5 .............1
4 x -3 y = 8 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 1
3 x 3 y = -15
4 x -3 y = 8
Add the two equations
7 x = -7
/ 7
x = -1
plug value of x in (1)
1 x + 1 y = -5
-1 + 1 y = -5
1 y = -5+1
1 y = -4

(-1,-4)
-2x+(1/2)y=1
multiply by 2
-4x+y=2..............(1)
2x+y=14..............(2)

-4 x + 1 y = 2 .............1
2 x + 1 y = 14 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
4 x -1 y = -2
2 x + 1 y = 14
Add the two equations
6 x = 12
/ 6
x = 2
plug value of x in (1)
-4 x + 1 y = 2
-8 + 1 y = 2
1 y = 2 + 8
1 y = 10
y = 10
(2,10)